OpenStudy (anonymous):

Can someone check over my work please? The area of a rectangular piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

OpenStudy (anonymous):

Here is my work: Part A: xy=A xy=280 x+1=y-5 x= y-6 Part B: (y-6)(y)=280 y^2-6y=280 y^2-6y-280=0 (y+14)(y-20) The width of the land is 20 meters.

OpenStudy (irishboy123):

have you thought about using "l" and "w" instead of "x" and "y" 'cos you may have got them the wrong way round.

OpenStudy (anonymous):

so length would be: l-5 and width would be: w+1 so does that mean I would have to set l-5=w+1?

OpenStudy (anonymous):

i dont think they're asking me to solve for x for part A, just to give them an equation...

OpenStudy (mrhoola):

I think I understand..

OpenStudy (mrhoola):

A rectangular area is usually denoted with 'L' for length and 'W' for width , since it is a rectangle L and W are not equal -Of course! but the statement then says that if the rectangular area is to be a square under two conditions , That is: i) L-5 ii) W +1 Right ? . So if we think about the square part , all sides are the same . since that is true we can use another variable , lets call it 'X' and we know the area of a square is x*x = Area

OpenStudy (anonymous):

so that would be my answer to part A? x= 280/(x+5)?

OpenStudy (mrhoola):

Do you understand what I did ??

OpenStudy (anonymous):

sort of?

OpenStudy (mrhoola):

x*x = area of a square however we were given a rectangle but they told us that if one side is L -5 and the other side W+1 would make that rectangular into a sqaure so I said , why not use the the area of the square equation X = L-5 and X = W +1 <--- This should totally make sense , right ?

OpenStudy (anonymous):

yes i think so??

OpenStudy (mrhoola):

area of a rectangle is L*W = area let L = X+5 so (x+5)*W = area finally W = area/(x+5)

OpenStudy (anonymous):

okay i understand that!

OpenStudy (anonymous):

is there something else i need to do for part a then?

OpenStudy (anonymous):

or is that it

OpenStudy (irishboy123):

@matenrouyes i am sorry if i have added to any confusion i was merely suggesting that you may have forgotten what x and y stand for when you started out, did x stand for length or width. i'd say the latter. so when you conclude rightly that y = 20 are you not concluding that length = 20 ?! that was the only point i was making and that was why i also suggested using more descriptive variable names

OpenStudy (irishboy123):

so i paste your entire post into a text editor, i switch w for x and l for y and i get this: Part A: wl=A wl=280 w+1=l-5 w= l-6 Part B: (l-6)(l)=280 l^2-6l=280 l^2-6l-280=0 (l+14)(l-20) see my point?

OpenStudy (anonymous):

OHHH I SEE!!

OpenStudy (irishboy123):

cool apart from that, it's excellent

OpenStudy (anonymous):

Thank you!! do you mind staying while i work it out to check it again?

OpenStudy (anonymous):

OK: so i think that for part a i had to find the equation of L, so this is what i did Part A: L-5=w+1 L= w+6 and then for part b, i would plug in what i got for A to get the width (w+6)(w)=280 and when i solve it i get -20, and 14, and obviously you cant have negative meters so the width is 14?? @IrishBoy123

OpenStudy (mrhoola):

oops sorry for the wrong answer . (u_u)

OpenStudy (irishboy123):

yes adding in a few of the details that an instructor might wish to see and that you already included above: \((w+6)w=280\) \(w^2 + 6w -280 = 0\) \((w+20)(w-14) = 0\) \(w = 14\) or \(w = -20\) so \(w = 14\)

OpenStudy (anonymous):

but can you agree that what i put for part A is correct?

OpenStudy (irishboy123):

ok i'd stick with what you did originally a part from the fact you forgot what x and y were representing so i pasted your work into a proessors and switched for w ad l and i got this: Part A: wl=A wl=280 w+1=l-5 w= l-6 that is what you originally wrote and i think it's great because it: 1) connects W and L in terms of the 280 area; and 2) connects W and L in terms of how they might be equal/ form a square

OpenStudy (irishboy123):

IOW, go with your original answer, just don't mix up width and length

OpenStudy (anonymous):

alright, thank you both so much! now can you tell me how to give you metals? lol

OpenStudy (irishboy123):

go up to a post by @Mrhoola and click Best Response

OpenStudy (anonymous):

okie dokie!! Thank you both <3

OpenStudy (irishboy123):

cool!!